OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10) and (-4, 0)? two fifths negative two fifths negative five halves five halves

OpenStudy (anonymous):

\[\Large Slope=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

slope = ((0-(-10))/(-4-0) =-10/4 =-5/2 negative five halves

OpenStudy (anonymous):

^right.

OpenStudy (anonymous):

-2.5

OpenStudy (anonymous):

negative five halves